What is the 10th term of a GP 5, 30, 180, .........
What is the sum of the first 5 terms?

Answer by richwmiller(17219)

Tn = t1 * r^(n - 1)
T10 = 5 * 10077696
T10 = 5 *10077696=50388480
T10 =50388480
Sn=t1*(1 - r^n)/(1 - r)
S5=5*(1 - 6^4)/(1 - 6)
S5=5*(1 - 1296)/(-5)
S5=-1*(-1295)
S5=1295